# Markarian 501 reincarnates with vigor as a temporary EHBL during VHE flaring in July 2014

### v1

@@ 4.1 Comparison with flaring events of 2005 and 2012
The spectral analysis of the VHE flaring events during July 16 to July 31, 2014 (total 15 days) shows that the flaring events in zone-1 were in high emission states like a standard HBL spectrum which we have studied previously. Also, the spectra in zone-1 are very similar to the flaring events of Mrk 501 observed during 2005 May-July and 2012 June (Sahu et al., 2020). Similar spectra were also observed from Mrk 421 in 2010 March (Sahu et al., 2021b). However, we observed a large variation in the VHE spectra of Mrk 501 in zone-2 during 2014 flaring event, particularly on July 19, 2014 (MAGIC Collaboration et al., 2020). On this day, the spectral index $\delta_{2}=4.0$ was observed to be very large, corresponding to a sharp fall in the VHE gamma-ray spectrum above a large cutoff energy $E^{c}_{\gamma}=3.18$ TeV.
During previous flaring events of Mrk 501 in 2005 and 2012, we did observe a variation in the spectra in zone-2, $3.1\lesssim\delta_{2}\lesssim 3.5$. However, this time the variation was observed to be abrupt. For example, on MJD 56856.91, the $E^{c}_{\gamma}=0.43$ TeV and $\delta_{2}=3.2$ and on next day MJD 56857.98 it jumped to 3.18 TeV and 4.0 respectively. Such large values were not seen before in any of the tEHBL flaring events. On this day, the intrinsic flux in zone-1 increased as $E^{0.17}_{\gamma}$ and decreased as $E^{-1}_{\gamma}$ in zone-2. Thus, at $E^{c}_{\gamma}=3.18$ TeV there is a peak-like structure in the intrinsic spectrum but manifests as a milder peak in the observed spectrum due to EBL smoothening effect. As the spectral indices $\beta_{1}$ and $\beta_{2}$ are the distribution of the background seed photons in zone-1 and zone-2 respectively, the tEHBL-like behavior of Mrk 501 must be attributed to SSC band where $\beta_{2}=2.0$ (zone-2). Previously, it was shown that the leptonic modeling of many HBL SEDs have deep valleys at the junction of synchrotron spectrum and the SSC spectrum with large $\beta$ value (Sahu et al., 2017, 2018, 2020, 2021b). It is possible that during these extreme flaring events, the inverse Compton scattering of high energy electrons with the self-produced synchrotron photons produces SSC photons in the narrow energy range of $3.1\,\mathrm{MeV}\lesssim\epsilon_{\gamma}\lesssim 6.6\,\mathrm{MeV}$ (zone-2), which are piled up and the flux behaves as $\Phi_{SSC}\propto\epsilon^{2}_{\gamma}$.
@@ 5 Discussion
Mrk 501 is a HBL which is extensively studied in all wavelenghts. Both leptonic models and hadronic models explain its spectra well. But the VHE emission during 2005, 2012 and 2014, when the synchrotron peak had shifted to frequencies above $10^{17}$ Hz, the VHE spectra were difficult to explain. Actually, during these flaring periods, Mrk 501 behaved like temporary EHBL and its spectra were complex. Similar tEHBL-like behavior were also observed from Mrk 421 and 1ES 1959+650. In previous studies, we used the two-zone photohadronic model and explained very well the tEHBL-like behavior of these sources. In the present work we have used the two-zone phtohadronic model again to explain the tEHBL-like behavior of Mrk 501 for the 15 days during July 16 to 31, 2014. It is observed that the spectra in zone-1 were like standard HBL and were in high emission states implying that the spectral indices in this region were in the range $3.02.6<\delta_{1}<$ 3.0. During the VHE flaring period of 2005 and 2012 in Mrk 501, this same behavior was observed. However, the spectral behavior in zone-2 of the faring in 2014 is much more complex and different then in the previous periods. Of particular importance is the VHE flaring event on MJD 56857.98 (July 19, 2014), when the cutoff energy $E^{c}_{\gamma}=3.18$ TeV and the spectral index in zone-2 was $\delta_{2}=4.0$. This high value of $\delta_{2}$ implies that the intrinsic spectrum falls as $F_{\gamma,in}\propto E^{-1}_{\gamma}$ for $E_{\gamma}>E^{c}_{\gamma}$. In zone-1, $F_{\gamma,in}\propto E^{0.17}_{\gamma}$, a slowly increasing flux up to $E^{c}_{\gamma}$. In the photohadronic model we predict a mild peak-like structure at $E^{c}_{\gamma}=3.18$ TeV in the observed spectrum. On this day, the maximum observed photon energy was $E_{\gamma}=6.8$ TeV, corresponding to a minimum bulk Lorentz factor $\Gamma\simeq 26.5$ which is consistent with the expectation from an EHBL. Except on 19th July, such spectral behavior was not observed on any other days. Also, such spectral behavior was not observed in the past from any of the HBLs, including Mrk 501.
@@
It is to be noted that our two-zone photohadronic model explains extremely well all the 15 spectra of the VHE flaring events of Mrk 501. Other nearby HBLs, which have already shown the temporary EHBL-like behavior in the past, may reincarnate again with vigor and behave like the flaring of July 19, 2014 of Mrk 501 or even more complex. However, many more flaring events from nearby HBLs have to be studied to establish this claim. With the present and/or future generation of IACTs it will be possible to monitor such flaring events in great detail to better understand their emission mechanisms and enrich theoretical modeling.
@@
The work of S.S. is partially supported by DGAPA-UNAM (Mexico) Project No. IN103522. A.U.P.O. and G. S.-C. would like to thank SECIHTI (Mexico) for partial support. Partial support from CSU-Long Beach is gratefully acknowledged. Pedro Fernández de Córdoba’s work is partially supported by the Spanish government through the RD projects PID2021-128676OB-I00 and PID2022-142407NB-I00, as well as by the Excellence Research Groups Program of the Generalitat Valenciana, Spain (PROMETEO CIPROM/2023/32).

### v2

@@ 4.1 Comparison with flaring events of 2005 and 2012
The spectral analysis of the VHE flaring events during July 16 to July 31, 2014 (total 15 days) shows that the flaring events in zone-1 were in high emission states like a standard HBL spectrum which we have studied previously. Also, the spectra in zone-1 are very similar to the flaring events of Mrk 501 observed during 2005 May-July and 2012 June Sahu et al. (2020). Similar spectra were also observed from Mrk 421 in 2010 March Sahu et al. (2021b). However, we observed a large variation in the VHE spectra of Mrk 501 in zone-2 during 2014 flaring event, particularly on July 19, 2014 MAGIC Collaboration and others (2020b). On this day, the spectral index $\delta_{2}=4.0$ was observed to be very large, corresponding to a sharp fall in the VHE gamma-ray spectrum above a large cutoff energy $E^{c}_{\gamma}=3.18$ TeV.
During previous flaring events of Mrk 501 in 2005 and 2012, we did observe a variation in the spectra in zone-2, $3.1\lesssim\delta_{2}\lesssim 3.5$. However, this time the variation was observed to be abrupt. For example, on MJD 56856.91, the $E^{c}_{\gamma}=0.43$ TeV and $\delta_{2}=3.2$ and on next day MJD 56857.98 it jumped to 3.18 TeV and 4.0 respectively. Such large values were not seen before in any of the tEHBL flaring events. On this day, the intrinsic flux in zone-1 increased as $E^{0.17}_{\gamma}$ and decreased as $E^{-1}_{\gamma}$ in zone-2. Thus, at $E^{c}_{\gamma}=3.18$ TeV there is a peak-like structure in the intrinsic spectrum but manifests as a milder peak in the observed spectrum due to EBL smoothening effect. As the spectral indices $\beta_{1}$ and $\beta_{2}$ are the distribution of the background seed photons in zone-1 and zone-2 respectively, the tEHBL-like behavior of Mrk 501 must be attributed to SSC band where $\beta_{2}=2.0$ (zone-2). We compare the values of $F_{2}$, $\delta_{2}$ and $E^{c}_{\gamma}$ of MJD 56857.98 with other days as summarized in Table 1 and plotted in Figs. 6 and 7. This comparison clearly shows that the spectrum of MJD 56857.98 is different from the rest.
Previously, it was shown that the leptonic modeling of many HBL SEDs have deep valleys at the junction of synchrotron spectrum and the SSC spectrum with large $\beta$ value Sahu et al. (2020, 2021a, 2021b, 2017, 2018a). It is possible that during these extreme flaring events, the inverse Compton scattering of high energy electrons with the self-produced background synchrotron photons produce SSC photons in the narrow energy range of $3.1\,\mathrm{MeV}\lesssim\epsilon_{\gamma}\lesssim 6.6\,\mathrm{MeV}$ (zone-2), which are piled up and the flux behaves as $\Phi_{SSC}\propto\epsilon^{2}_{\gamma}$ and for $\epsilon_{\gamma}\gtrsim 6.6\,\mathrm{MeV}$ (zone-1) the flux is flatter behaving as $\Phi_{SSC}\propto\epsilon^{0.83}_{\gamma}$. Thus, the high energy protons interacting with these background SSC photons will produce the observed VHE gamma-ray spectrum.
@@ 5 Discussion
Mrk 501 is a HBL which is extensively studied in all wavelengths. Both leptonic models and hadronic models explain its spectra well. But the VHE emission during 2005, 2012 and 2014, when the synchrotron peak had shifted to frequencies above $10^{17}$ Hz, the VHE spectra were difficult to explain. Actually, during these flaring periods, Mrk 501 behaved like tEHBL and its spectra were complex. Similar tEHBL-like behavior also observed from Mrk 421 and 1ES 1959+650. In previous studies, we used the two-zone photohadronic model and explained very well the tEHBL-like behavior of these sources. In the present work we have used the two-zone photohadronic model again to explain the tEHBL-like behavior of Mrk 501 for the 15 days during July 16 to 31, 2014. It is observed that the spectra in zone-1 were like standard HBL and were in high emission states implying that the spectral indices in this region were in the range $2.6<\delta_{1}<3.0$. During the VHE flaring period of 2005 and 2012 in Mrk 501, this same behavior was observed. However, the spectral behavior in zone-2 of the flaring in 2014 is much more complex and different than in the previous periods. Of particular importance is the VHE flaring event on MJD 56857.98 (July 19, 2014), when the cutoff energy $E^{c}_{\gamma}=3.18$ TeV and the spectral index in zone-2 was $\delta_{2}=4.0$. This high value of $\delta_{2}$ implies that the intrinsic spectrum falls as $F_{\gamma,in}\propto E^{-1}_{\gamma}$ for $E_{\gamma}>E^{c}_{\gamma}$. In zone-1, $F_{\gamma,in}\propto E^{0.17}_{\gamma}$, a slowly increasing flux up to $E^{c}_{\gamma}$. In the photohadronic model we predict a mild peak-like structure at $E^{c}_{\gamma}=3.18$ TeV in the observed spectrum. On this day, the maximum observed photon energy was $E_{\gamma}=6.8$ TeV, corresponding to a minimum bulk Lorentz factor $\Gamma\simeq 26.5$ which is consistent with the expectation from an EHBL. Except on 19th July, such spectral behavior was not observed on any other days. Also, such spectral behavior was not observed in the past from any of the HBLs, including Mrk 501. It is expected that there may also be narrow TeV feature in the VHE spectrum on July 18 (one day before) and/or on July 20 (one day after). However, the photohadronic model does not predict any such feature on these days.
In the context of the two-zone photohadronic model, the VHE flaring event of MJD 56857.98 is due to the interaction of high energy protons with the low energy SSC background photons. In zone-1, protons of energy $E_{p}\lesssim 31.8\,\mathrm{TeV}$ interact with the SSC photons of energy $\epsilon_{\gamma}\gtrsim\,6.6\,\mathrm{MeV}$ to produce the observed VHE gamma-ray spectrum of energy $E_{\gamma}\lesssim 3.18\,\mathrm{TeV}$. In zone-2, protons in the energy range $31.8\,\mathrm{TeV}\lesssim E_{p}\lesssim 68\,\mathrm{TeV}$ interact with the SSC photons in the energy range $3.1\,\mathrm{MeV}\lesssim\epsilon_{\gamma}\lesssim 6.6\,\mathrm{MeV}$ to produce the VHE gamma-ray spectrum in the energy range $3.18\,\mathrm{TeV}\lesssim E_{\gamma}\lesssim 6.8\,\mathrm{TeV}$. The VHE spectra of other days are similar to the flaring event of MJD 56857.98 but have different spectral behavior of the SSC photons in zone-1 and zone-2 and also have different cutoff energy $E^{c}_{\gamma}$.
On MJD 56865.99 and MJD 56868.01, the spectra are relatively smooth which could probably be fitted with a single power-law. However, during the whole observation period, the synchrotron peak of Mrk 501 was always above $10^{17}$ Hz, a clear signature of transition to an EHBL. Thus, the VHE spectra must be dealt with two-zone photohadronic model Sahu et al. (2020, 2021a, 2021b). On the contrary, by using a single power-law, there will be no consistent way of dealing with these spectra. The two-zone approach is therefore not an arbitrary choice, but a necessary condition to reproduce the spectral shape and their temporal evolution during different flaring periods when the synchrotron peak is above $10^{17}$ Hz.
@@
It is to be noted that our two-zone photohadronic model explains well all the 15 spectra of the VHE flaring events of Mrk 501. Other nearby HBLs, which have already shown the tEHBL-like behavior in the past, may reincarnate again with vigor and behave like the flaring of July 19, 2014 of Mrk 501 or even more complex. However, many more flaring events from nearby HBLs have to be studied to establish this claim. With the present and/or future generation of IACTs it will be possible to monitor such flaring events in great detail to better understand their emission mechanisms and enrich theoretical modeling.
Acknowledgements.
@@
The work of S.S. is partially supported by DGAPA-UNAM (Mexico) Project No. IN105326. A.U.P.O., D.I.P.S and G. S.-C. would like to thank SECIHTI (Mexico) for partial support. Partial support from CSU-Long Beach is gratefully acknowledged. Pedro Fernández de Córdoba’s work is partially supported by the Spanish government through the RD projects PID2021-128676OB-I00 and PID2022-142407NB-I00, as well as by the Excellence Research Groups Program of the Generalitat Valenciana, Spain (PROMETEO CIPROM / 2023 / 32).